### v1

# FEDONet : Fourier-Embedded DeepONet for Spectrally Accurate Operator Learning

Abstract
Deep Operator Networks (DeepONets) have recently emerged as powerful data-driven frameworks for learning nonlinear operators, particularly suited for approximating solutions to partial differential equations (PDEs). Despite their promising capabilities, the standard implementation of DeepONets, which typically employs fully connected linear layers in the trunk network, can encounter limitations in capturing complex spatial structures inherent to various PDEs. To address this, we introduce Fourier-embedded trunk networks within the DeepONet architecture, leveraging random Fourier feature mappings to enrich spatial representation capabilities. Our proposed Fourier-embedded DeepONet, FEDONet demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the two-dimensional Poisson equation, Burgers’ equation, the Lorenz-63 chaotic system, Eikonal equation, Allen–Cahn equation, Kuramoto–Sivashinsky equation, and the Lorenz-96 system. Empirical evaluations of FEDONet consistently show significant improvements in solution reconstruction accuracy, with average relative $\ell_{2}$ performance gains ranging between 2-3x compared to the DeepONet baseline. This study highlights the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
1 Introduction
Partial differential equations (PDEs) underpin models ranging from turbulent-flow dynamics and heat diffusion to electromagnetics and biological transport. Classical numerical solvers—finite-difference, finite-element and spectral methods—provide high-fidelity solutions but incur a prohibitive computational burden when faced with high-dimensional, multiscale or stiff systems [1, 2, 3]. The quest for rapid surrogates has a long pedigree: universal-approximation theorems for neural networks [4, 5, 6, 7], reduced-order techniques based on proper-orthogonal decomposition (POD) [8, 9], kernel–Gaussian-process surrogates [10], and mesh-free radial-basis schemes [11, 12]. Yet, translating these ideas into *function-to-function* (operator) mappings that respect arbitrary discretisations remained elusive.
The modern wave of operator learning addresses this gap by training neural networks that act directly on infinite-dimensional function spaces and generalise across grid resolutions. Deep Operator Networks (DeepONets) [13] and the broader Neural Operator family [14, 15, 16] establish a branch–trunk architecture with rigorous operator–universal-approximation guarantees. Fourier Neural Operators (FNOs) extend this paradigm through global spectral convolutions, delivering impressive speed-ups on parametric PDEs [17]. Subsequent variants exploit multiwavelets [18, 19], convolutional kernels [20], multi-grid tensorisation [21, 22] and geometry-aware deformations [23] to boost resolution or handle complex domains.
Beyond kernel choices, researchers have tackled discontinuities [24], derivative supervision [25], irregular meshes via graph trunks [26], active multi-resolution sampling [27], and U-Net/wavelet hybrids [28, 29]. Random Fourier features have proved especially effective at encoding high-frequency structure in standard deep nets [30], yet remain under-explored inside operator networks whose trunk is typically a plain fully connected multilayer perceptron (MLP).
To incorporate governing laws and uncertainty, Physics-Informed Neural Networks (PINNs) [31, 32] evolved into Physics-Informed Neural Operators (PINO) [33, 34] and their variational or pseudo-physics relatives [35, 36]. Complementary efforts embed operators in latent or invertible spaces for data efficiency and bidirectional inference [37, 38, 39, 40].
Transformer-style attention has recently been adopted for operator learning [41, 42, 43, 44, 45] and for geometry generalisation [46]. Such architectures underpin landmark climate-scale surrogates such as FourCastNet [47] and its amortised FNO successor [48].
Despite the rich taxonomy above, the trunk of conventional DeepONets, charged with encoding spatial/temporal coordinates, remains a shallow MLP. This design struggles to capture the oscillatory or multi-scale patterns ubiquitous in nonlinear PDE solutions, leading to sub-optimal sample efficiency and accuracy. Recent trunk alternatives (e.g. graph [26] or wavelet U-Net [28]) introduce architectural complexity or domain-specific preprocessing.
We propose a Fourier-Embedded DeepONet (FEDONet) that replaces the raw coordinates fed into the trunk with a compact random Fourier-feature map [30]. The embedding supplies an implicit spectrum of sinusoidal basis functions, endowing the trunk with a richer hypothesis space at negligible parameter cost. FEDONet preserves the original branch–trunk separation, is plug-and-play with existing DeepONet codebases, and adds no runtime overhead beyond a single matrix multiplication.
  1. (i)
    We formulate the Fourier embedding of trunk coordinates, derive its effect on the operator neural tangent kernel, and show that it strictly enlarges the approximation class of DeepONets.
  2. (ii)
    We deliver a thorough benchmark across seven PDE families—steady Poisson, viscous Burgers, Allen–Cahn, Kuramoto–Sivashinsky, Eikonal, Lorenz–63 and Lorenz–96—demonstrating 48.93% average reduction in relative $\ell^{2}$ error versus a tuned vanilla DeepONet.
  3. (iii)
    We analyse hyperparameter sensitivity (mapping size and scale) and empirically confirm that the Fourier trunk improves convergence speed without sacrificing discretisation invariance.
By injecting a principled spectral bias into the DeepONet trunk, FEDONet closes the performance gap to frequency-domain operators such as FNO while retaining the flexibility and locality advantages of the original architecture.
2 Methodology
We aim to develop a data-driven framework that approximates nonlinear operators mapping between infinite-dimensional function spaces. Let $\Omega\subset\mathbb{R}^{D}$ be a bounded domain, and define the input and output function spaces:
$\mathcal{U}$ $=\left\{u:\mathcal{X}\rightarrow\mathbb{R}^{d_{u}}\right\},\quad\mathcal{X}\subseteq\mathbb{R}^{d_{x}},$ (1)
$\mathcal{S}$ $=\left\{s:\mathcal{Y}\rightarrow\mathbb{R}^{d_{s}}\right\},\quad\mathcal{Y}\subseteq\mathbb{R}^{d_{y}},$ (2)
where $\mathcal{U}$ denotes input functions (e.g., boundary or initial conditions) and $\mathcal{S}$ represents output fields (e.g., PDE solutions). Our goal is to learn an approximation $\mathcal{G}_{\theta}:\mathcal{U}\rightarrow\mathcal{S}$ to an unknown operator $\mathcal{G}$, where $\theta\in\Theta$ are trainable parameters.
In practice, we are given a dataset $\mathcal{D}=\{(u^{i},s^{i})\}_{i=1}^{N}$, comprising function pairs sampled from $\mathcal{G}$. DeepONet is a neural operator architecture well-suited to this task. It decomposes the learning problem into two components: Branch network $B_{\theta}$ that encodes discrete samples of $u$ at sensor points $\{x_{1},\dots,x_{m}\}$, Trunk network $T_{\theta}$ that takes in spatial (or spatiotemporal) coordinates $\zeta=(x,y,z,t)$.
The predicted value at location $\zeta$ is computed via an inner product:
$\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\zeta).$ (3)
2.1 Fourier Embeddings as Spectral Preconditioners
Standard MLPs exhibit a well-known spectral bias — they learn low-frequency components faster than high-frequency ones [49]. This bottleneck is particularly harmful in operator learning problems involving sharp gradients, multiscale dynamics, or oscillatory fields.
To address this, we propose injecting fixed Fourier embeddings into the trunk network input. Given a coordinate $\zeta\in\mathbb{R}^{d}$, we define a randomized Fourier feature map:
$\phi(\zeta)=\left[\sin(2\pi B\zeta),\cos(2\pi B\zeta)\right],\quad B_{ij}\sim\mathcal{N}(0,\sigma^{2}),$ (4)
where $B\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ — enriching its spectral content.
The DeepONet prediction then becomes:
$\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\phi(\zeta))=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)).$ (5)
Figure 1: FEDONet: Fourier Embedded Deep operator Network
This transformation can be interpreted as a randomized kernel approximation:
$k(\zeta,\zeta^{\prime})\approx\phi(\zeta)^{\top}\phi(\zeta^{\prime}),$ (6)
effectively lifting $\zeta$ into a high-dimensional basis of sinusoidal functions. By Bochner’s theorem [50], such embeddings approximate shift-invariant kernels — implicitly kernelizing the trunk network. This expands its representational capacity, enabling better learning of high-frequency targets.
Beyond expressivity, Fourier embeddings improve the conditioning of the learning process. Under mild assumptions, the embedded features exhibit approximate whitening ( see A ) :
$\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I,$ (7)
reducing feature correlations and leading to a more isotropic Neural Tangent Kernel (NTK) spectrum. This in turn improves optimization dynamics, stabilizes gradient flow, and accelerates convergence during training.
Let $\mathcal{H}_{\phi}$ denote the function class learnable with Fourier features, and $\mathcal{H}_{\text{vanilla}}$ the class learnable with raw coordinates. Then:
$\mathcal{H}_{\phi}\supset\mathcal{H}_{\text{vanilla}},$ (8)
with the inclusion being strict in problems with rich frequency content (see B). Fourier embeddings thus act as spectral preconditioners that align the input geometry with the task’s spectral demands.
2.2 Training Objective
We adopt a supervised learning framework to train the proposed FEDONet model. The objective is to approximate a target operator $G$ by minimizing the prediction error over a dataset of paired input–output functions. Each dataset element consists of a function $u^{(i)}\in\mathcal{U}$ and corresponding evaluations of the output $s^{(i)}_{j}=G(u^{(i)})(\zeta^{(i)}_{j})$ at query locations $\zeta^{(i)}_{j}\in\mathcal{Y}$.
Let the dataset be denoted by
$\mathcal{D}=\left\{\left(u^{(i)},\left\{\zeta^{(i)}_{j},s^{(i)}_{j}\right\}_{j=1}^{Q}\right)\right\}_{i=1}^{N},$
where $N$ is the number of input functions and $Q$ is the number of query points per function.
The model prediction at location $\zeta$ is given by:
$G_{\theta}(u)(\zeta)=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)),$
where $b_{k}(u)$ and $t_{k}(\phi(\zeta))$ denote the outputs of the branch and Fourier-embedded trunk networks, respectively.
The training objective is to minimize the empirical risk using the mean squared error (MSE) loss:
$\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{Q}\sum_{j=1}^{Q}\left\|G_{\theta}(u^{(i)})(\zeta^{(i)}_{j})-s^{(i)}_{j}\right\|^{2}.$
Model parameters are optimized using mini-batch stochastic gradient descent with the Adam optimizer. Learning rate schedules are selected based on the problem characteristics to ensure stable and efficient convergence. Each mini-batch consists of randomly sampled function–query pairs $(u,\zeta)$, enabling scalable and efficient training across diverse domains.
Notably, the Fourier embedding layer $\phi(\zeta)$ is fixed and non-trainable throughout training. This design choice ensures low overhead while enhancing model expressivity through spectral enrichment. Nonetheless, the architecture can be extended to incorporate learnable embeddings $\phi_{\theta}(\zeta)$, enabling adaptive spectral tuning tailored to problem-specific frequency content.
2.3 Evaluation and Generalization
After training, the FEDONet model is evaluated on unseen input functions $u_{\text{test}}$. To assess model performance, we report both quantitative and qualitative metrics:
  • Relative $L^{2}$ Error: Defined as
    $\varepsilon_{L^{2}}=\frac{\|G_{\theta}(u)-G(u)\|_{2}}{\|G(u)\|_{2}},$
    where the norm is computed over a dense evaluation grid. This metric quantifies the normalized discrepancy between predicted and reference output fields.
  • Spectral Fidelity: We perform Fourier analysis of the predicted and true fields to evaluate preservation of high-frequency content. Specifically, we compare the angle-integrated power spectra $E(k)$ across wavenumbers $k$, assessing the model’s ability to capture multiscale dynamics and resolve fine-scale features.
The ability to evaluate the learned operator at arbitrary coordinates makes FEDONet particularly suitable for scientific computing tasks requiring continuous surrogate modeling. In all benchmark scenarios, FEDONet demonstrates strong generalization, enhanced spectral accuracy, and robust performance across a wide range of PDE systems with varying stiffness, nonlinearity, and complexity.
3 Results
3.1 2D Poisson Equation
We created a synthetic dataset comprising numerical solutions of the two-dimensional (2D) Poisson equation, a fundamental PDE commonly employed in modeling steady-state processes such as heat conduction, electrostatics, and diffusion phenomena. The mathematical formulation of the Poisson equation is given by:
$\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in[0,1]^{2},$ (9)
accompanied by homogeneous Dirichlet boundary conditions:
$u(x,y)\big{|}_{\partial\Omega}=0.$ (10)
In this context, $u(x,y)$ denotes the solution field, while $f(x,y)$ represents a specified source term or coefficient.
To generate the dataset, we synthesized $N=10,000$ distinct instances of the coefficient function $f$ over a uniform spatial grid of resolution $128\times 128$. Each instance of the coefficient field was produced using Gaussian Random Fields (GRFs) characterized by specific spectral parameters, notably $\alpha=3$ and $\tau=3$, resulting in:
$f\sim\text{GRF}(\alpha=3,\tau=3).$ (11)
Once these coefficient fields were created, we numerically solved the associated Poisson equation to obtain the corresponding solution fields $u(x,y)$. The numerical discretization employed a finite difference approach, utilizing a conventional five-point stencil approximation of the Laplacian operator. By discretizing the spatial domain $\Omega=[0,1]^{2}$ into a $128\times 128$ structured grid, the PDE was transformed into a linear algebraic system:
$A\mathbf{u}=b,$ (12)
where $\mathbf{u}\in\mathbb{R}^{16384}$ denotes the vectorized solution field and $b\in\mathbb{R}^{16384}$ is the vector representation of the source term $f$. The discretized Laplacian matrix $A$ is defined explicitly as:
[…]
with homogeneous Dirichlet boundary conditions enforced directly by setting boundary rows of $A$ as identity matrices and corresponding entries in $f$ to zero.
Ultimately, this methodology yielded 10,000 independent pairs of coefficient-source fields and their respective solutions, forming a robust dataset suitable for training, validating, and benchmarking neural operator models. The dataset facilitates a comprehensive evaluation of the trade-offs in computational efficiency, accuracy, and reconstruction performance of our proposed neural operator architectures.
Figure 2 presents a representative test sample from the two–dimensional Poisson dataset. The input forcing field is shown on the left, followed by the exact solution, the prediction of a vanilla DeepONet, and the corresponding point-wise error. The bottom row repeats the comparison for our Fourier-Embedded DeepONet (FEDONet).
  • Spatial reconstruction. FEDONet reproduces the lobed extrema, smooth gradients and corner behaviour of the ground truth almost indistinguishably, whereas the vanilla network exhibits visible blurring and slight spatial shifts of the high-magnitude regions. The residual of the vanilla model oscillates between $\pm 1.5\times 10^{-3}$, displaying high-frequency artefacts; FEDONet reduces this range to $\pm 3\times 10^{-4}$, a five-fold improvement that manifests as a low-amplitude, coherent pattern.
  • Energy spectra. Figure 3 plots the angle-integrated energy spectrum $E_{k}$ for four randomly chosen hold-out fields (samples 29, 15, 24 and 17). Both networks match the reference spectrum at large scales ($k\lesssim 5$), but qualitative differences arise in the inertial and dissipation ranges:
    1. (a)
      The vanilla DeepONet (purple) consistently over-estimates energy for $k\gtrsim 8$, producing a spectral “pile-up” that grows by nearly an order of magnitude in sample 24—mirroring the spurious small-scale oscillations seen in its error map.
    2. (b)
      FEDONet (red) tracks the ground truth (dark blue) almost perfectly up to the Nyquist limit, deviating by at most a factor of two in the far dissipation tail and thereby preserving the correct energy cascade.
Figure 2: Model reconstruction comparison for the 2D Poisson Dataset.
Together, the spatial and spectral diagnostics confirm that Fourier embeddings in the trunk endow DeepONets with the high-frequency resolution and stability required for accurate operator learning: FEDONet not only slashes point-wise errors but also eliminates the artificial energy plateau that plagues fully-connected trunks.
Figure 3: Energy spectrum comparison between the ground truth, vanilla and Fourier DeepOnet reconstructions for the 2D Poisson Dataset.
3.2 Burgers’ Equation
The Burgers’ equation is a well-known 1D viscous equation. It has been studied extensively for DeepONets in prior literature. The equation is given by:
$\frac{\partial u}{\partial t}(x,t)+u\frac{\partial u}{\partial x}(x,t)=\nu\frac{\partial^{2}u}{\partial x^{2}}(x,t),\>\>\>\>\>\forall(x,t)\in[0,1]\times[0,1]$ (14)
where $x$ and $t$ denote the spatio-temporal coordinates and $\nu$ is the kinematic viscosity. We set $\nu=0.01$ along with initial and periodic boundary conditions as:
$\begin{split}\displaystyle u(x,0)=&\>s(x),\>\>\>\>\>\>\>\>\>\>\>\forall\>x\in[0,1]\\ u(0,t)=&\>u(1,t),\>\>\>\>\>\>\>\>\forall\>t\in[0,1]\\ \frac{\partial u}{\partial x}(0,t)=&\>\frac{\partial u}{\partial x}(1,t)\>\>\>\>\>\>\forall\>t\in[0,1]\end{split}$ (15)
where the initial condition $s(x)$ is sampled from a Gaussian Random Field, $s(x)\sim\mathcal{N}\left(0,25^{2}(-\Delta+5^{2}I)^{-4}\right)$, satisfying periodicity. The objective is to learn the nonlinear operator $\mathcal{G}$ mapping $s(x)$ to the full spatio-temporal solution $u(x,t)$ using both vanilla DeepONet and the proposed FEDONet.
Figure 4 shows the relative $\ell_{2}$ errors over 128 unseen test samples. FEDONet consistently outperforms DeepONet, with a mean error reduction from 4.01% to 2.63%, and a narrower spread in error distribution indicating increased reliability.
Figure 4: Burgers Relative $\ell_{2}$ error across 128 unseen samples for DeepONet and FEDONet.
Figure 5 presents the best-performing test sample, showing the ground truth solution alongside predictions from both models. While DeepONet captures the coarse solution structure, it struggles with steep gradients and exhibits visible smoothing near shock regions. In contrast, FEDONet produces significantly sharper and more accurate reconstructions, faithfully capturing shock fronts and fine-scale dynamics.
Figure 5: Best-performing test sample: Ground truth vs. DeepONet vs. FEDONet prediction.
The energy spectra of the predicted and ground truth fields, shown in Figure 6, further emphasize this advantage. FEDONet accurately captures the spectral decay across a wide range of wavenumbers, aligning with the physical expectations of dissipative transport phenomena. DeepONet, however, underestimates energy at higher modes, reflecting its inability to resolve fine features and spectral richness. This demonstrates that the Fourier embeddings effectively bridge the model’s spectral bias, enabling it to better represent high-frequency dynamics essential to Burgers-type flows.
These results underscore the importance of aligning model inductive bias with the physical spectrum of the target PDE. The enhanced performance of FEDONet in both physical and spectral domains highlights its potential as a more robust and generalizable operator learner for convection-dominated systems.
Figure 6: Temporal Evolution of Pointwise Error and Energy Spectrum: Best-Performing Sample vs. Ground Truth
Figure 7 shows the convergence of training losses for different batch sizes. FEDONet consistently converges faster and to lower loss values across all training configurations, indicating improved optimization characteristics and stability during training.
Figure 7: Training loss convergence for the 1D Burgers’ Equation: Comparison of Vanilla and Fourier Embedded DeepONets trained with varying batch sizes.
3.3 Lorenz–63 System
The Lorenz–63 system is a widely studied chaotic dynamical system that serves as a benchmark for evaluating the long-term stability and predictive accuracy of surrogate models. It is governed by the following system of coupled nonlinear ordinary differential equations:
$\frac{dx}{dt}$ $=\sigma(y-x),$ (16)
$\frac{dy}{dt}$ $=x(\rho-z)-y,$
$\frac{dz}{dt}$ $=xy-\beta z,$
where the canonical parameters are set to $\sigma=10$, $\rho=28$, and $\beta=8/3$. Under these settings, the system exhibits chaotic dynamics on a strange attractor in three-dimensional phase space.
The objective is to approximate the solution operator $\mathcal{G}\colon x_{0}\mapsto\mathbf{x}(t)$, where $\mathbf{x}(t)=(x(t),y(t),z(t))$ denotes the full trajectory over the time interval $t\in[0,3]$, given a scalar initial condition $x_{0}\in[10,15]$, with $y_{0}=z_{0}=12$ held fixed. This configuration reduces the input to a one-dimensional function-to-function mapping, while preserving the rich, sensitive behavior of the output trajectory.
A dataset of 10,00010{,}000 trajectories is generated using the classic Runge–Kutta method (RK45), each evaluated at 1000 uniformly spaced time steps. The input to the DeepONet branch network is the scalar $x_{0}$, and the trunk network is queried at time coordinates $t\in[0,3]$. The goal is to accurately recover the full state trajectory, including lobe transitions and fine-scale oscillations characteristic of chaotic dynamics.
Figure 8: Reconstructed Lorenz–63 attractor for a randomnly sampled initial condition
Figure 8 shows the reconstructed attractor for a representative test sample. While DeepONet departs from the ground truth early in the rollout, FEDONet maintains geometric coherence and captures both wings of the attractor. This result demonstrates the stabilizing effect of spectral enrichment on trajectory learning.
Figure 9: Phase-plane projection of the xxzz states. The vanilla DeepONet trajectory diverges from the expected orbit, showing geometric drift and incorrect lobe transitions. FEDONet reconstructs the spiral arms and inter-lobe dynamics with higher accuracy, preserving the overall topology.
In Figure 9, we examine the xxzz projection. The DeepONet prediction deviates from the true phase-space trajectory and fails to capture the correct lobe transitions. In contrast, FEDONet successfully recovers the attractor’s phase geometry, indicating more robust learning of long-term dependencies.
Figure 10: Pointwise prediction error in $x(t)$, $y(t)$, and $z(t)$ for a representative sample. The vanilla DeepONet exhibits compounding errors, especially after $t>1.5$, reflecting the accumulation of approximation drift. FEDONet achieves significantly lower errors and better error containment beyond lobe transitions, highlighting improved stability under chaotic evolution.
Figure 10 presents the pointwise prediction errors over time. DeepONet accumulates noticeable errors, particularly in the $x(t)$ and $z(t)$ components, which undergo rapid transitions between lobes. These deviations grow sharply after $t>1.5$, illustrating compounding approximation errors. FEDONet exhibits substantially lower and smoother error profiles, with better robustness during nonlinear state transitions.
Figure 11: Cumulative relative error over time. DeepONet exhibits rapid exponential error growth beyond $t>1.5$, characteristic of chaotic divergence. FEDONet accumulates error more gradually, indicating improved long-term stability and mitigation of runaway prediction drift.
In Figure 11, we observe that the cumulative error of DeepONet grows exponentially, characteristic of chaotic divergence. In contrast, FEDONet maintains a flatter cumulative error curve, demonstrating enhanced resilience against long-term instability and better preservation of trajectory fidelity.
Figure 12: Spectral energy density of predicted $x(t)$ trajectories compared with the ground truth. DeepONet underestimates energy in mid-to-high frequency bands, resulting in oversmoothing and loss of chaotic detail. FEDONet preserves the spectral tail and matches the broadband structure more accurately, contributing to improved phase and amplitude resolution.
Finally, Figure 12 compares the spectral content of the predicted and reference trajectories. The vanilla DeepONet fails to capture high-frequency modes, resulting in a smoothed trajectory that diverges from the correct dynamics. FEDONet, equipped with randomized Fourier features, preserves the broadband energy spectrum more accurately—an essential factor in recovering the fine-scale structure and temporal richness of chaotic systems.
These results collectively demonstrate that FEDONet offers improved phase alignment, lower pointwise and cumulative error, and superior spectral accuracy in chaotic dynamical systems. The introduction of Fourier embeddings into the trunk network acts as a spectral preconditioner, enabling the model to overcome inductive biases and improve fidelity under sensitivity to initial conditions.
3.4 Eikonal Equation
We consider the two-dimensional Eikonal equation, which governs the propagation of wavefronts at unit speed and arises in a wide range of applications including geometric optics, level-set methods, and computational geometry. The equation is given by:
$\|\nabla s(\mathbf{x})\|_{2}$ $=1,$ (17)
$s(\mathbf{x})$ $=0,\quad\mathbf{x}\in\partial\Omega,$
where $\mathbf{x}=(x,y)\in\mathbb{R}^{2}$ denotes the spatial coordinates, $\Omega\subset\mathbb{R}^{2}$ is the computational domain, and $\partial\Omega$ is its boundary. The solution $s(\mathbf{x})$ represents the signed distance from a point $\mathbf{x}\in\Omega$ to the boundary $\partial\Omega$, defined as:
[…]
with the distance function $d$ defined by:
$d(\mathbf{x},\partial\Omega):=\inf_{\mathbf{y}\in\partial\Omega}\|\mathbf{x}-\mathbf{y}\|_{2}.$ (19)
In this study, we aim to learn the solution operator that maps a binary mask of a two-dimensional geometry to its corresponding signed distance field (SDF). As a testbed, we construct a dataset of NACA 4-digit airfoil shapes, widely used in aerodynamics, and compute their SDFs within a structured computational domain.
A NACA 4-digit airfoil is parameterized by its maximum camber $m$, the location of maximum camber $p$, and the maximum thickness $t$. The mean camber line is given by:
$y_{c}(x)=\begin{cases}\frac{m}{p^{2}}(2px-x^{2}),&0\leq x<p,\\ \frac{m}{(1-p)^{2}}((1-2p)+2px-x^{2}),&p\leq x\leq 1,\end{cases}$ (20)
and its derivative is:
$\frac{dy_{c}}{dx}=\begin{cases}\frac{2m}{p^{2}}(p-x),&0\leq x<p,\\ \frac{2m}{(1-p)^{2}}(p-x),&p\leq x\leq 1.\end{cases}$ (21)
The thickness distribution $y_{t}(x)$ is defined as:
$y_{t}(x)=5t\left(0.2969\sqrt{x}-0.1260x-0.3516x^{2}+0.2843x^{3}-0.1036x^{4}\right).$ (22)
The upper and lower surface coordinates of the airfoil are computed as:
$x_{u}=x-y_{t}\sin(\theta),\quad$ $y_{u}=y_{c}+y_{t}\cos(\theta),$ (23)
$x_{l}=x+y_{t}\sin(\theta),\quad$ $y_{l}=y_{c}-y_{t}\cos(\theta),$ (24)
where $\theta=\tan^{-1}\left(\frac{dy_{c}}{dx}\right)$ is the local inclination angle of the camber line.
Each airfoil is embedded in a $256\times 256$ pixel computational grid and scaled to fit within a central subdomain. A binary mask is created to define the geometry, and the signed distance function is computed using the Euclidean Distance Transform (EDT) as:
$s(x,y)=d_{\text{out}}(x,y)-d_{\text{in}}(x,y),$ (25)
where $d_{\text{out}}$ and $d_{\text{in}}$ represent the unsigned distances outside and inside the airfoil boundary, respectively. To ensure numerical stability, the resulting SDF is normalized to the range [1,1][-1,1]:
$s_{\text{norm}}(x,y)=\frac{s(x,y)}{\max|s(x,y)|}.$ (26)
A total of 1250 distinct airfoil shapes are synthesized by randomly sampling parameters $m\in[0.01,0.09]$, $p\in[0.1,0.7]$, and $t\in[0.1,0.4]$. This dataset captures a wide range of geometrical variations and curvature profiles, providing a rigorous benchmark for learning PDE solution operators from geometric input.
Figure 13: Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry. Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
As shown in Figure 13, FEDONet achieves more accurate recovery of the level-set structure. The DeepONet underperforms in regions of high curvature, particularly near the airfoil’s leading and trailing edges, where geometric non-smoothness introduces steep spatial gradients. FEDONet’s use of Fourier embeddings in the trunk network enhances its ability to resolve these features by expanding the function space with oscillatory basis functions aligned with the data’s spectral demands.
These results confirm that FEDONet offers improved resolution of high-gradient and discontinuous regions in PDE fields, even in geometry-conditioned tasks such as the Eikonal equation. This highlights its potential as a general-purpose surrogate for PDEs defined on complex, spatially varying domains.
3.5 Lorenz-96 Chaotic System
To evaluate the performance of our DeepONet-based neural operator models, we utilize the Lorenz-96 (L96) system—a well-known chaotic dynamical system commonly employed for benchmarking data-driven models and surrogate learning techniques in climate modeling and nonlinear dynamics. The L96 system is defined by the set of coupled ordinary differential equations:
$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,\quad i=1,2,\dots,N$ (27)
with periodic boundary conditions: $x_{i-N}=x_{i}=x_{i+N}$. Here, $x_{i}$ denotes the state of the $i$-th variable, $N$ is the number of state variables, and $F$ is a constant forcing parameter that controls the nonlinearity and chaotic behavior of the system. For $F>8$, the system is known to exhibit strong chaos. In our experiments, we select $F=4.0$, which yields nonlinear but non-chaotic dynamics.
The L96 system is integrated forward in time using the classical fourth-order Runge-Kutta (RK4) method. Given an initial condition $\mathbf{x}_{0}\in\mathbb{R}^{N}$, the RK4 method computes the solution at the next time step using:
$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4})$ (28)
where the intermediate slopes $k_{1},k_{2},k_{3},k_{4}$ are computed using the Lorenz-96 time derivative evaluated at successive RK stages, and $\Delta t=0.01$ is the time step.
To construct a comprehensive dataset for training and evaluating neural operator models, we generate an ensemble of 10,00010{,}000 independent trajectories of the L96 system:
  • The system dimension is set to $N=40$.
  • Initial conditions are generated by perturbing the equilibrium state $x_{i}=F$ with small Gaussian noise:
    $\mathbf{x}_{0}=F\cdot\mathbf{1}+\epsilon\cdot\mathcal{N}(0,I),\quad\epsilon=10^{-3}$
  • Each trajectory is simulated for a total time of 15 seconds.
  • To remove transient behavior and capture steady-state dynamics, the first 10 seconds (burn-in period) are discarded.
  • The remaining 5 seconds of simulation are retained, corresponding to $500+1=501$ time steps with $\Delta t=0.01$.
The resulting dataset has shape (10,000,501,40)(10{,}000,501,40), where each sample corresponds to a trajectory of 40-dimensional state vectors over 501 time steps. This dataset captures the spatiotemporal evolution of the system from a diverse set of initial conditions and serves as a challenging testbed for learning nonlinear operator mappings.
Figure 14: Complete spatio-temporal field reconstruction comparison between DeepONet and FEDONet.
Figure 14 displays two representative hold-out trajectories of the L96 system. For each sample we plot the initial state, the reference spatio-temporal evolution, predictions from the DeepONet and FEDONet, together with the corresponding absolute-error fields.
  • Phase propagation. Both networks reproduce the diagonal stripe pattern that characterises L96 wave propagation, preserving phase speed and wavelength over the 500-step horizon. No visible drift or amplitude blow-up appears in either solution.
  • Error magnitude. The residual maps reveal only modest differences: in Sample 1 the RMS error of the vanilla model is $\mathcal{O}(10^{-1})$ while FEDONet trims it to the lower end of the same decade; Sample 2 shows the opposite ordering at isolated time–state patches. Overall, FEDONet yields a slightly lower global $\ell_{2}$ error but the gap is far narrower than that observed for the Poisson or Burgers benchmarks.
  • Interpretation. Because the L96 solution manifold is dominated by a single, slowly varying spatio-temporal mode, a shallow fully connected trunk already provides a sufficiently rich basis. Injecting random Fourier features therefore offers limited extra benefit, consistent with the quantitative numbers in Table 1, where both models attain sub-25 % relative error.
3.6 Allen–Cahn Equation
We evaluate the performance of the proposed Fourier-Embedded DeepONet (FEDONet) in modeling the one-dimensional Allen–Cahn equation—a nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems. The governing partial differential equation is given by:
$\frac{\partial u}{\partial t}=\epsilon\frac{\partial^{2}u}{\partial x^{2}}-5u^{3}+5u,\quad x\in[-1,1],\quad t\in[0,1],$ (29)
where $u(x,t)$ is the phase field variable and $\epsilon=10^{-4}$ denotes the diffusion coefficient. Periodic boundary conditions are enforced on both $u$ and its spatial derivative $\partial u/\partial x$, consistent with the physical assumption of a closed domain.
To construct a challenging dataset with rich spatial features, we generate 10,000 solution trajectories using an explicit Euler solver. The spatial and temporal grids are discretized using $\Delta x=0.01$ and $\Delta t=0.005$, yielding solution arrays of dimension $200\times 200$. Initial conditions are defined as:
$s(x)=\sum_{k=1}^{3}\left[a_{k}x^{2k}\cos(k\pi x)+b_{k}x^{2k}\sin(k\pi x)\right],$ (30)
where $a_{k},b_{k}\sim\mathcal{U}(0,1)$ are independent samples from the uniform distribution. This construction introduces multiscale structure and localized steep gradients into the initial data, posing a nontrivial test for surrogate learning.
Each training sample comprises a pair $(s(x),u(x,t))$, where $s(x)\in\mathbb{R}^{200\times 1}$ denotes the initial profile, and $u(x,t)\in\mathbb{R}^{200\times 200}$ is the full spatiotemporal solution. The goal is to approximate the solution operator $\mathcal{G}\colon s(x)\mapsto u(x,t)$, comparing the performance of vanilla DeepONet and FEDONet.
Figure 15: Best-performing test case for the Allen–Cahn equation. Relative $\ell_{2}$ error: $4.85\%$ for DeepONet, $2.27\%$ for FEDONet.
Figure 15 shows qualitative predictions for the best-performing test case. FEDONet sharply resolves the interface zones and phase transitions, while DeepONet produces smoother approximations with notable blurring near discontinuities. The spectral inductive bias introduced by randomized Fourier features enables FEDONet to retain high-frequency details crucial to phase-field dynamics.
Figure 16: Energy Spectrum and Pointwise absolute error across space and time for a representative sample.
The pointwise absolute error plot in Figure 16 emphasizes the localized improvements of FEDONet in capturing steep interfaces. DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
Figure 16 compares the spatial energy spectra of predicted and true solutions. FEDONet accurately tracks the ground truth spectrum across all wavenumbers, preserving the high-frequency tail associated with sharp phase boundaries. In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen–Cahn system.
These results collectively demonstrate that the incorporation of spectral priors via Fourier embeddings in FEDONet enhances both pointwise accuracy and spectral fidelity. The improved resolution of phase interfaces and better retention of high-frequency information position FEDONet as a robust surrogate model for nonlinear, multiscale PDEs such as the Allen–Cahn equation.
3.7 Kuramoto–Sivashinsky Equation
We evaluate the performance of FEDONet on the one-dimensional Kuramoto–Sivashinsky (KS) equation, a canonical model that exhibits spatiotemporal chaos due to the combined effects of nonlinear advection, destabilizing diffusion, and stabilizing fourth-order hyperviscosity. The governing equation is given by:
$\begin{split}&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{4}u}{\partial x^{4}}=0,\quad x\in[0,L],\;t\in[0,T],\\ &u(x,0)=u_{0}(x),\quad x\in[0,L],\end{split}$ (31)
where $u(x,t)$ is the scalar field of interest, and $L=24$, $T=50$ denote the spatial and temporal domains respectively. The initial condition $u_{0}(x)$ is synthesized as a randomized Fourier series:
$u_{0}(x)=\sum_{n=1}^{4}C_{n}\sin\left(\frac{nx}{L}\right),\quad C_{n}\sim\mathcal{N}(0,1),$ (32)
which introduces low-frequency coherent structures that quickly evolve into chaotic waveforms under the KS dynamics.
We generate 10,000 such initial conditions and simulate the corresponding solution fields using a pseudo-spectral solver. The objective is to learn the operator $\mathcal{G}\colon u_{0}(x)\mapsto u(x,t)$ using both the standard DeepONet and the proposed FEDONet models. This problem serves as a stringent benchmark due to the sensitivity of chaotic systems to initial conditions and their broadband frequency content.
Figure 17: Reconstruction of a median-performing KS trajectory. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
Figure 17 shows the reconstructed spatiotemporal field for a median test case. The FEDONet model effectively captures the nonlinear evolution of the solution, preserving the wave envelope, phase dynamics, and intermittent modulation typical of KS attractors. In stark contrast, the vanilla DeepONet produces incoherent and smoothed reconstructions, with pronounced phase lag and energy collapse beyond $t>20$. This degradation is reflected in a high relative error of $89.63\%$ compared to just $16.36\%$ for FEDONet.
In Figure 18, we highlight the best-performing test case. FEDONet achieves a relative $L^{2}$ error of only $10.77\%$, accurately modeling both the fine-scale chaotic filaments and large-scale rolling structures. Notably, the trajectory remains consistent with the ground truth over long integration times, demonstrating that the spectral trunk can stabilize predictions and suppress drift. Such performance underscores the model’s ability to retain energy in high-frequency modes—essential for modeling non-periodic fine-grained behavior that typifies chaotic fields.
Figure 18: Best-performing KS test sample. FEDONet retains the amplitude and frequency content of the chaotic trajectory, yielding a low relative $L^{2}$ error of $10.77\%$.
Figure 19: Worst-performing test sample. Even in failure cases, FEDONet remains qualitatively reasonable. Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
Figure 19 displays the worst-case test result. While the FEDONet prediction exhibits growing deviations from the reference field, it still retains discernible wave structures and phase coherence, with an error of $58.02\%$. In comparison, the vanilla DeepONet fails entirely, producing numerical artifacts, amplitude distortion, and full trajectory collapse. This robustness under extreme conditions emphasizes the value of frequency-enhanced representations in learning dynamical system operators.
Figure 20 plots the training loss curves for both models. FEDONet converges faster and exhibits more stable optimization behavior, attributable to the improved conditioning of the input space via randomized Fourier features. This aligns with observations from kernel learning and feature engineering literature, where frequency-based transformations are known to facilitate smoother descent landscapes and gradient propagation.
Figure 20: Training convergence comparison. FEDONet achieves faster and more stable optimization due to implicit spectral regularization.
Figure 21: Distribution of relative $L^{2}$ errors across 512 unseen test samples. FEDONet maintains lower error in both mean and variance, demonstrating improved robustness and generalization in chaotic regimes.
Finally, Figure 21 summarizes performance statistics across 512 randomly selected test cases. FEDONet consistently achieves lower relative $L^{2}$ errors with reduced variance, indicating better generalization and resilience to initial condition perturbations.
In totality, the Kuramoto–Sivashinsky benchmark clearly illustrates the limitations of standard DeepONet trunks in handling chaotic, broadband dynamics. The introduction of Fourier embeddings into the trunk network provides an effective inductive bias that promotes spectral alignment, temporal coherence, and long-horizon accuracy. FEDONet emerges as a compelling model for operator learning in complex regimes, particularly those characterized by sensitive dependence on initial conditions and multiscale feature interactions.
4 Summary
In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations (PDEs). By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training—all without increasing runtime complexity. FEDONet consistently outperforms the vanilla DeepONet. Notably, it delivers dramatic reductions in relative $\ell_{2}$ error—most prominently in multiscale and high-frequency scenarios such as the Kuramoto–Sivashinsky equation (76% $\rightarrow$ 21%).
Table 1: Relative $\ell_{2}$ error (%) for each benchmark.
Dataset DeepONet FEDONet
2-D Poisson (elliptic) 5.80 2.32
1-D Burgers (conv./diff.) 4.01 2.63
Lorenz–63 (chaotic ODE) 2.03 0.46
Lorenz–96 (chaotic ODE) 23.40 22.90
Allen–Cahn (phase–field) 10.61 5.50
Kuramoto–Sivashinsky (spatio-chaos) 75.11 20.62
Eikonal (Hamilton–Jacobi) 2.15 1.12
Table 2: Qualitative performance across seven benchmark problems.
Method 2-D Poisson 1-D Burgers Lorenz–63 Lorenz–96 Allen–Cahn KS Eikonal
DeepONet
FEDONet ✓✓ ✓✓ ✓✓ ✓✓
  • Legend: ✓ decent;  ✓✓ excellent;  ✗ sub-par.
Beyond pointwise accuracy, FEDONet preserves physical consistency across modalities. Energy spectrum analyses reveal that FEDONet better captures inertial and dissipation range behavior, avoiding the artificial spectral pile-up observed in DeepONets. In geometry-sensitive tasks like the Eikonal equation, it reconstructs sharp features and discontinuities with higher fidelity.
Taken together, these results establish FEDONet as a robust, scalable, and architecture-agnostic enhancement for neural operator learning. The proposed spectral trunk serves as a drop-in replacement that aligns the inductive biases of DeepONets with the spectral demands of complex PDE systems, providing a promising path forward for efficient surrogate modeling in scientific machine learning.
5 Future Work
The Fourier-Embedded DeepONet (FEDONet) architecture presents a compelling foundation for enhancing operator learning, yet several promising extensions remain unexplored. A natural next step is to generalize beyond sinusoidal embeddings by investigating alternative basis functions in the trunk network. Wavelets, Chebyshev polynomials, or orthogonal polynomials tailored to specific PDE families could offer improved localization or boundary adaptivity, especially for systems exhibiting sharp interfaces or multi-scale dynamics. Additionally, hybridizing spectral embeddings with structured architectures such as graph-based trunks or encoder–decoder networks may further boost performance on irregular domains or hierarchical solution spaces.
Another avenue involves making the Fourier embedding parameters learnable. While the current design uses fixed random frequency matrices, enabling adaptive tuning of these frequencies could allow the model to better align with the intrinsic spectral content of the data. This modification could yield a more expressive and data-efficient hypothesis space while preserving the training stability benefits of approximate whitening. Finally, incorporating uncertainty quantification and physics-informed constraints into the FEDONet framework—via Bayesian formulations or PINO-style residual losses—may further enhance robustness, interpretability, and generalization in scientific computing tasks involving sparse or noisy data.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This work was supported in part by the AFOSR Grant FA9550-24-1-0327.
Data availability
Data supporting the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Whitening Effect of Fourier Feature Embeddings
Let $\phi(\zeta)=\sqrt{2}\,[\sin(2\pi B\zeta),\cos(2\pi B\zeta)]\in\mathbb{R}^{2M}$ be the Fourier feature embedding of an input $\zeta\in\mathbb{R}^{d}$, where each row of the matrix $B\in\mathbb{R}^{M\times d}$ is sampled i.i.d. from a Gaussian distribution:
$B_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (33)
Assume further that the input $\zeta$ is uniformly distributed over a compact domain, e.g., $\zeta\sim\mathcal{U}([0,1]^{d})$. Under these assumptions, we show that the random Fourier features $\phi(\zeta)$ are approximately whitened, i.e.,
$\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I_{2M}.$ (34)
To see this, consider a single frequency row vector $B_{i}\in\mathbb{R}^{1\times d}$ drawn from $B$. The corresponding $2$-dimensional block of the embedding is:
$\phi_{i}(\zeta)=\sqrt{2}\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)\\ \cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}$. (35)
We now compute the second-order statistics of this embedding block under the assumption that $B_{i}\cdot\zeta$ is uniformly distributed over $[0,1]$ (a reasonable approximation due to the randomness in $B_{i}$ and the uniformity of $\zeta$). Using standard trigonometric integrals, we obtain:
$\mathbb{E}_{\zeta}[\sin^{2}(2\pi B_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (36)
$\mathbb{E}_{\zeta}[\cos^{2}(2\pi B_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (37)
$\mathbb{E}_{\zeta}[\sin(2\pi B_{i}\cdot\zeta)\cos(2\pi B_{i}\cdot\zeta)]$ $\approx 0.$ (38)
Therefore, the unscaled covariance matrix for this block is approximately:
$\mathbb{E}_{\zeta}\left[\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)\\ \cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)&\cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}\right]\approx\frac{1}{2}I_{2}.$ (39)
Scaling by the $\sqrt{2}$ factor yields:
$\mathbb{E}_{\zeta}[\phi_{i}(\zeta)\phi_{i}(\zeta)^{\top}]\approx I_{2}.$ (40)
Since the rows of $B$ are sampled independently and the embeddings for each $B_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies:
$\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I_{2M}.$ (41)
Hence, the random Fourier features are approximately whitened. This whitening effect implies that the feature coordinates are uncorrelated and have unit variance. In practice, this leads to improved numerical conditioning and convergence in gradient-based optimization, particularly via the Neural Tangent Kernel (NTK), which benefits from isotropic input distributions. As such, Fourier embeddings not only enhance expressivity but also promote well-conditioned training dynamics.
Appendix B Theoretical Justification for Fourier Embedding Superset Property
In this appendix, we formally justify the claim stated in Section 2.1 that the hypothesis space induced by Fourier embeddings strictly contains the vanilla hypothesis space of coordinate-input networks. We present this as a lemma and provide constructive proof using classical function approximation theory.
Let $\mathcal{H}_{\text{vanilla}}$ be the hypothesis space of functions representable by a feedforward neural network with input $\zeta\in\mathbb{R}^{d}$, and let $\mathcal{H}_{\phi}$ be the hypothesis space of functions representable by a similar network where the input is first mapped via the Fourier feature embedding:
$\phi(\zeta)=\left[\sin(2\pi B\zeta),\cos(2\pi B\zeta)\right],\quad B\in\mathbb{R}^{M\times d},\;B_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (42)
Then, under mild regularity assumptions (e.g., compact input domain $K\subset\mathbb{R}^{d}$ and continuous activation functions), we have:
$\mathcal{H}_{\text{vanilla}}\subsetneq\mathcal{H}_{\phi}.$ (43)
We divide the proof into two parts.
Part I: Inclusion. Let $f\in\mathcal{H}_{\text{vanilla}}$ be a function representable by a neural network of the form:
$f(\zeta)=\sum_{j=1}^{N}a_{j}\,\sigma(w_{j}^{\top}\zeta+b_{j}).$ (44)
Since each $\zeta_{i}$ (coordinate function) can be approximated by a Fourier series on a compact domain $K$, and the Fourier embedding $\phi(\zeta)$ contains these sinusoidal components, it follows that a function of $\zeta$ can be emulated by a function of $\phi(\zeta)$. In particular, for every $f\in\mathcal{H}_{\text{vanilla}}$ and $\varepsilon>0$, there exists $\tilde{f}\in\mathcal{H}_{\phi}$ such that:
$\sup_{\zeta\in K}|f(\zeta)-\tilde{f}(\zeta)|<\varepsilon.$ (45)
Therefore, $\mathcal{H}_{\text{vanilla}}\subseteq\mathcal{H}_{\phi}$.
Part II: Strict Containment. Consider the function $f(\zeta)=\sin(100\pi\zeta)$ over $\zeta\in[0,1]$. Due to the spectral bias of neural networks [49], this high-frequency function is difficult to approximate using shallow networks on raw input. However, if the frequency matrix $B$ used in the embedding $\phi(\zeta)$ includes rows close to $100$, then:
$\sin(2\pi B_{j}\zeta)\approx\sin(100\pi\zeta),$
making $f$ easily representable in $\mathcal{H}_{\phi}$.
Since $f(\zeta)$ cannot be efficiently approximated in $\mathcal{H}_{\text{vanilla}}$ (requiring significantly more parameters or depth), but is trivially accessible in $\mathcal{H}_{\phi}$, we conclude that the inclusion is strict:
$\mathcal{H}_{\text{vanilla}}\subsetneq\mathcal{H}_{\phi}.$ (46)
This result formalizes the intuition that Fourier embeddings expand the expressivity of the network and can be interpreted as a form of spectral preconditioning for operator learning models such as DeepONets.
Appendix C Spectral Interpretation of Fourier-Embedded DeepONets
The use of Fourier embeddings in the trunk network induces a natural spectral decomposition in the DeepONet output. Specifically, the composition of the trunk and branch networks can be interpreted as a data-driven spectral projection:
$\mathcal{G}(u)(\zeta)\approx\sum_{k=1}^{p}\langle u,\varphi_{k}\rangle\,\psi_{k}(\zeta),$ (47)
where $\{\varphi_{k}\}_{k=1}^{p}$ are data-dependent basis functions learned from the branch network, and $\{\psi_{k}(\zeta)\}_{k=1}^{p}$ are basis functions defined by the trunk network, shaped by the sinusoidal feature embedding $\phi(\zeta)$.
When the trunk network is fed sinusoidal embeddings of the form
$\phi(\zeta)=[\cos(2\pi\omega_{1}^{\top}\zeta),\dots,\cos(2\pi\omega_{M}^{\top}\zeta),\sin(2\pi\omega_{1}^{\top}\zeta),\dots,\sin(2\pi\omega_{M}^{\top}\zeta)],$
the basis functions $\psi_{k}(\zeta)$ are learned combinations of trigonometric modes. This allows the model to represent oscillatory and localized patterns that are characteristic of multiscale physical systems.
The representation in Equation (47) mirrors classical Galerkin methods, where the solution is projected onto a known orthonormal basis:
$u(x,t)\approx\sum_{n=1}^{N}a_{n}(t)\,\phi_{n}(x),$ (48)
with the coefficients $a_{n}(t)$ governed by PDE dynamics. In DeepONet, both the basis functions $\psi_{k}$ and the projection coefficients $\langle u,\varphi_{k}\rangle$ are learned directly from data, enabling generalization beyond handcrafted bases.
Unlike classical global bases such as Fourier or Chebyshev functions, the neural trunk network can generate localized, adaptive basis functions. This enhances the model’s ability to approximate functions with sharp transitions, discontinuities, or compactly supported features, which are common in real-world PDE systems. We may formalize the DeepONet architecture as a composition of a projection and synthesis operator:
$\Pi_{\varphi}(u):=[\langle u,\varphi_{1}\rangle,\dots,\langle u,\varphi_{p}\rangle]^{\top},\quad\mathcal{S}_{\psi}(\zeta):=[\psi_{1}(\zeta),\dots,\psi_{p}(\zeta)]^{\top},$
yielding the output:
$\mathcal{G}(u)(\zeta)\approx\mathcal{S}_{\psi}(\zeta)^{\top}\Pi_{\varphi}(u),$
which resembles a spectral synthesis operation of the form:
$f(x)\approx\sum_{k}\hat{f}_{k}\phi_{k}(x).$
This interpretation leads to several insights:
  1. 1.
    Fourier-embedded DeepONets naturally align with the dominant spectral modes of the target operator, facilitating more efficient approximation of high-frequency content.
  2. 2.
    The learned basis can adapt to complex patterns in the data, unlike fixed basis expansions.
  3. 3.
    The operator learning framework parallels well-established numerical discretization techniques, such as Petrov–Galerkin and spectral methods, but operates in a fully data-driven regime.
In summary, Fourier embeddings serve as a principled mechanism for spectral lifting, enabling DeepONet to perform a learned Galerkin-style decomposition of operators. This perspective bridges classical PDE approximation theory with modern deep learning, offering both interpretability and flexibility in learning complex function-to-function mappings.