### v3

# FEDONet: Fourier-Embedded DeepONet for Spectrally Accurate Operator Learning

Abstract
Deep Operator Networks (DeepONets) have recently emerged as powerful data-driven frameworks for learning nonlinear operators, particularly suited for approximating solutions to partial differential equations. Despite their promising capabilities, the standard implementation of DeepONets, which typically employs fully connected linear layers in the trunk network, can encounter limitations in capturing complex spatial structures inherent to various PDEs. To address this limitation, we introduce Fourier-Embedded trunk networks within the DeepONet architecture, leveraging random fourier feature mappings to enrich spatial representation capabilities. Our proposed Fourier-Embedded DeepONet, FEDONet demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the two-dimensional Poisson, Burgers’, Lorenz-63, Eikonal, Allen-Cahn, and the Kuramoto-Sivashinsky equation. FEDONet delivers consistently superior reconstruction accuracy across all benchmark PDEs, with particularly large relative $L^{2}$ error reductions observed in chaotic and stiff systems. This study highlights the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
keywords:
Neural Operator Learning , Fourier Feature Embeddings , Deep Operator Networks , Partial Differential Equations
journal: Journal of Computational Physics
[…]
organization=Mechanical and Aerospace Engineering, addressline=1512 Middle Drive, University of Tennessee, city=Knoxville, postcode=37996-2210, state=TN, country=USA
1 Introduction
Partial differential equations underpin models ranging from turbulent‐flow dynamics and heat diffusion to electromagnetics and biological transport. Classical numerical solvers such as finite‐difference, finite‐element and spectral methods provide high‐fidelity solutions but incur a prohibitive computational burden when faced with high‐dimensional, multiscale or stiff systems [1, 2, 3]. Efforts to construct efficient surrogate models have a long history, spanning the universal approximation theorems for neural networks[4, 5, 6, 7], reduced‐order techniques based on proper‐orthogonal decomposition [8, 9], kernel Gaussian process surrogates [10], and mesh‐free radial‐basis schemes [11, 12]. Early data‐driven approaches also sought to learn governing equations directly from trajectory data using deep neural networks [13] However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long standing challenge.
The recent emergence of operator learning addresses this long-standing challenge by training neural networks that act directly on infinite-dimensional function spaces, enabling generalization across discretizations and grid resolutions. Deep Operator Networks [14] pioneered the branch-trunk architecture for learning nonlinear operators, while subsequent Neural Operator formulations [15, 16, 17] extended these ideas with rigorous operator-approximation guarantees and scalable architectures. Building on these foundations, Fourier Neural Operators [18] employ global spectral convolutions to capture nonlocal correlations efficiently, achieving remarkable speed-ups and accuracy gains in parametric PDE learning. Related ideas have been explored in the context of learning PDE dynamics in modal or Fourier space [19], where the evolution operator is approximated directly on spectral coefficients rather than physical fields, highlighting the efficacy of frequency-space representations for capturing multiscale dynamics. Subsequent developments have further advanced this paradigm through multiwavelet representations [20, 21], convolutional kernels [22], multigrid tensorizations [23, 24], and geometry-aware spectral deformations [25], collectively enhancing scalability, resolution fidelity, and geometric adaptability.
Beyond architectural and kernel innovations, recent research has expanded operator learning along multiple frontiers. Efforts addressing solution irregularities have focused on handling discontinuities [26], incorporating derivative supervision for enhanced gradient fidelity [27], and enabling learning on irregular meshes through graph-based trunk networks [28]. Adaptive and multi-resolution strategies have been introduced to improve sampling efficiency [29], while hybrid encoder-decoder frameworks leveraging U-Net and wavelet representations further enhance multi-scale expressivity [30, 31]. To embed physical constraints and quantify uncertainty, the paradigm of Physics-Informed Neural Networks (PINNs) [32, 33] has been generalized into Physics-Informed Neural Operators [34, 35] and related variational or pseudo-physics extensions [36, 37]. Parallel efforts have explored latent and invertible operator formulations to achieve data-efficient and bidirectional inference [38, 39, 40, 41]. Transformer based attention mechanisms have also been adapted for operator learning [42, 43, 44, 45, 46] and for robust geometry generalization [47]. Together, these advances have enabled high-fidelity, climate-scale surrogates such as FourCastNet [48] and its amortized Fourier Neural Operator successor [49], underscoring the maturity and scalability of the operator-learning paradigm.
Despite the recent operator-learning architectures, the trunk network of conventional DeepONets remains a shallow multilayer perceptron. This design often struggles to represent the oscillatory and multi-scale structures that characterize nonlinear PDE solutions, resulting in reduced sample efficiency and reconstruction accuracy. While recent alternatives such as graph-based [28] or wavelet-enhanced U-Net trunks [30] offer improved expressivity, they introduce significant architectural complexity and domain-specific preprocessing requirements. In contrast, random Fourier features have proven remarkably effective at encoding high-frequency information in standard deep networks [50], yet their potential within operator-learning frameworks remains largely unexplored, where the trunk is typically implemented as a simple fully connected MLP. To address this gap, we propose the Fourier-Embedded Deep Operator Network (FEDONet), which augments the DeepONet trunk by replacing raw spatial coordinates with a compact random Fourier-feature embedding [50]. This embedding implicitly introduces a spectrum of sinusoidal basis functions, enriching the trunk’s representational capacity with negligible additional parameters. FEDONet retains the original branch-trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
We formally introduce the Fourier embedding of trunk coordinates and analyze its theoretical implications through the lens of the operator neural tangent kernel, demonstrating that the embedding strictly enlarges the approximation class of standard DeepONets. A comprehensive benchmark is carried spanning seven canonical PDE families - steady Poisson, viscous Burgers, Allen-Cahn, Kuramoto-Sivashinsky, Eikonal, and Lorenz-63. By embedding a principled spectral bias into the trunk network, FEDONet bridges the performance gap to frequency-domain operators, while preserving the architectural simplicity, flexibility, and locality benefits of the original DeepONet formulation. Section  2 presents the formulation of the Fourier-Embedded Deep Operator Network (FEDONet), detailing the embedding construction, its theoretical connection to the operator neural tangent kernel, and its implications for approximation capacity. Section  3 describes the benchmark setup and implementation details across the seven PDE families considered. Sections  4 and  5 summarize the main findings and outlines directions for future research in spectral operator learning and physics-constrained network design.
2 Methodology
We aim to develop a data-driven framework that approximates nonlinear operators mapping between infinite-dimensional function spaces. Let $\Omega\subset\mathbb{R}^{D}$ be a bounded domain, and define the input and output function spaces:
$\mathcal{U}$ $=\left\{u:\mathcal{X}\rightarrow\mathbb{R}^{d_{u}}\right\},\quad\mathcal{X}\subseteq\mathbb{R}^{d_{x}},$ (1)
$\mathcal{S}$ $=\left\{s:\mathcal{Y}\rightarrow\mathbb{R}^{d_{s}}\right\},\quad\mathcal{Y}\subseteq\mathbb{R}^{d_{y}},$ (2)
where $\mathcal{U}$ denotes input functions (e.g., boundary or initial conditions) and $\mathcal{S}$ represents output fields (e.g., PDE solutions). Our goal is to learn an approximation $\mathcal{G}_{\theta}:\mathcal{U}\rightarrow\mathcal{S}$ to an unknown operator $\mathcal{G}$, where $\theta\in\Theta$ are trainable parameters.
In practice, we are given a dataset $\mathcal{D}=\{(u^{i},s^{i})\}_{i=1}^{N}$, comprising function pairs sampled from $\mathcal{G}$. DeepONet is a neural operator architecture well-suited to this task. It decomposes the learning problem into two components: Branch network $B_{\theta}$ that encodes discrete samples of $u$ at sensor points $\{x_{1},\dots,x_{m}\}$, Trunk network $T_{\theta}$ that takes in spatial (or spatiotemporal) coordinates $\zeta=(x,y,z,t)$.
The predicted value at location $\zeta$ is computed via an inner product:
$\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\zeta).$ (3)
2.1 Fourier Embeddings as Spectral Preconditioners
Standard MLPs exhibit a well-known spectral bias - they learn low-frequency components faster than high-frequency ones [51]. This bottleneck is particularly harmful in operator learning problems involving sharp gradients, multiscale dynamics, or oscillatory fields.
To address this, we propose injecting fixed Fourier Embeddings into the trunk network input illustrated in Fig. 1. Given a coordinate $\zeta\in\mathbb{R}^{d}$, we define a randomized Fourier feature map:
$\phi(\zeta)=\left[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)\right],\quad Z_{ij}\sim\mathcal{N}(0,\sigma^{2}),$ (4)
where $Z\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ - enriching its spectral content.
The FEDONet prediction then becomes:
$\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\phi(\zeta))=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)).$ (5)
Figure 1: FEDONet: Fourier Embedded Deep operator Network
This transformation can be interpreted as a randomized kernel approximation:
$k(\zeta,\zeta^{\prime})\approx\phi(\zeta)^{\top}\phi(\zeta^{\prime}),$ (6)
effectively lifting $\zeta$ into a high-dimensional basis of sinusoidal functions. By Bochner’s theorem [52], such embeddings approximate shift-invariant kernels - implicitly kernelizing the trunk network. This expands its representational capacity, enabling better learning of high-frequency targets.
Beyond expressivity, Fourier embeddings improve the conditioning of the learning process. Under mild assumptions, the embedded features exhibit approximate whitening ( see A ) :
$\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I,$ (7)
reducing feature correlations and leading to a more isotropic Neural Tangent Kernel (NTK) spectrum. This in turn improves optimization dynamics, stabilizes gradient flow, and accelerates convergence during training.
Let $\mathcal{H}_{\phi}$ denote the function class learnable with Fourier features, and $\mathcal{H}_{\text{vanilla}}$ the class learnable with raw coordinates. Then:
$\mathcal{H}_{\phi}\supset\mathcal{H}_{\text{vanilla}},$ (8)
with the inclusion being strict in problems with rich frequency content (see B). Fourier embeddings thus act as spectral preconditioners that align the input geometry with the task’s spectral demands.
2.2 Training Objective
We adopt a supervised learning framework to train the proposed FEDONet model. The objective is to approximate a target operator $G$ by minimizing the prediction error over a dataset of paired input-output functions. Each dataset element consists of a function $u^{(i)}\in\mathcal{U}$ and corresponding evaluations of the output $s^{(i)}_{j}=G(u^{(i)})(\zeta^{(i)}_{j})$ at query locations $\zeta^{(i)}_{j}\in\mathcal{Y}$.
Let the dataset be denoted by
$\mathcal{D}=\left\{\left(u^{(i)},\left\{\zeta^{(i)}_{j},s^{(i)}_{j}\right\}_{j=1}^{Q}\right)\right\}_{i=1}^{N},$
where $N$ is the number of input functions and $Q$ is the number of query points per function.
The model prediction at location $\zeta$ is given by:
$G_{\theta}(u)(\zeta)=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)),$
where $b_{k}(u)$ and $t_{k}(\phi(\zeta))$ denote the outputs of the branch and Fourier-embedded trunk networks, respectively.
The training objective is to minimize the empirical risk using the mean squared error (MSE) loss:
$\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{Q}\sum_{j=1}^{Q}\left\|G_{\theta}(u^{(i)})(\zeta^{(i)}_{j})-s^{(i)}_{j}\right\|^{2}.$
Model parameters are optimized using mini-batch stochastic gradient descent with the Adam optimizer. Learning rate schedules are selected based on the problem characteristics to ensure stable and efficient convergence. Each mini-batch consists of randomly sampled function-query pairs $(u,\zeta)$, enabling scalable and efficient training across diverse domains.
Notably, the Fourier embedding layer $\phi(\zeta)$ is fixed and non-trainable throughout training. This design choice ensures low overhead while enhancing model expressivity through spectral enrichment. Nonetheless, the architecture can be extended to incorporate learnable embeddings $\phi_{\theta}(\zeta)$, enabling adaptive spectral tuning tailored to problem-specific frequency content.
2.3 Evaluation and Generalization
After training, the FEDONet model is evaluated on unseen input functions $u_{\text{test}}$. To assess model performance, we report both quantitative and qualitative metrics:
  • Relative $L^{2}$ Error: Defined as
    $\varepsilon_{L^{2}}=\frac{\|G_{\theta}(u)-G(u)\|_{2}}{\|G(u)\|_{2}},$
    where the norm is computed over a dense evaluation grid. This metric quantifies the normalized discrepancy between predicted and reference output fields.
  • Spectral Fidelity: We perform Fourier analysis of the predicted and true fields to evaluate preservation of high-frequency content. Specifically, we compare the angle-integrated power spectra $E(k)$ across wavenumbers $k$, assessing the model’s ability to capture multiscale dynamics and resolve fine-scale features.
The ability to evaluate the learned operator at arbitrary coordinates makes FEDONet particularly suitable for scientific computing tasks requiring continuous surrogate modeling. In all benchmark scenarios, FEDONet demonstrates strong generalization, enhanced spectral accuracy, and robust performance across a wide range of PDE systems with varying stiffness, nonlinearity, and complexity.
3 Results
3.1 2D Poisson Equation
We constructed a synthetic dataset comprising numerical solutions to the canonical two-dimensional Poisson equation. The governing equation is
$\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in[0,1]^{2},$ (9)
subject to homogeneous Dirichlet boundary conditions,
$u(x,y)\big|_{\partial\Omega}=0,$ (10)
where $u(x,y)$ denotes the potential field and $f(x,y)$ represents a prescribed source distribution.
A total of N=10, [^0] N=10{,}000 distinct source fields were synthesized using Gaussian random fields (GRFs) with spectral parameters $\alpha=3$ and $\tau=3$,
$f\sim\text{GRF}(\alpha=3,\tau=3),$ (11)
on a uniform $128\times 128$ grid. For each realization of $f$, the corresponding solution $u$ was obtained by numerically solving the Poisson equation using a standard five-point finite-difference discretization of the Laplacian operator. The spatial domain $\Omega=[0,1]^{2}$ was thus transformed into a structured algebraic system,
$A\mathbf{u}=\mathbf{b},$ (12)
where $\mathbf{u},\mathbf{b}\in\mathbb{R}^{16384}$ denote the vectorized forms of the solution and source fields, respectively. The coefficient matrix $A$ was defined using a conventional five-point stencil with Dirichlet conditions enforced by setting boundary rows to the identity and zeroing the corresponding entries in $\mathbf{b}$. This process yielded 10,000 pairs of input-output fields suitable for training, validation, and testing of neural operator architectures.
Figure 2: Solving a 2D Poisson equation: Comparison of predicted and reference solution fields of DeepONet and FEDONet
Figure 2 shows a representative test case from the Poisson dataset. The top row depicts the input forcing field, exact solution, prediction from the baseline DeepONet, and the corresponding pointwise error, while the bottom row presents results from the proposed Fourier-Embedded DeepONet (FEDONet). Both models were evaluated on $1000$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $1.41\%$ for the DeepONet and $1.09\%$ for the FEDONet. FEDONet reproduces the reference field with sharper gradients and more accurate spatial localization of extrema. The residual field of the vanilla DeepONet displays fine-scale artifacts, while FEDONet yields a smoother, spatially coherent error distribution, corresponding to an approximate fivefold improvement in reconstruction accuracy.
Figure 3: Energy spectrum comparison between the ground truth, vanilla and Fourier DeepOnet reconstructions for the 2D Poisson Dataset.
Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ shown in Fig. 3. Both models reproduce the reference spectrum at large scales ($k\lesssim 50$), but differences arise in the intermediate and dissipative ranges. The baseline DeepONet overestimates energy for $k\gtrsim 90$, resulting in a spurious buildup of high-wavenumber content that mirrors the oscillatory residuals seen in physical space. In contrast, FEDONet closely follows the reference spectrum across the entire range. Taken together, the spatial and spectral analyses demonstrate that the Fourier embedding enhances the trunk network’s ability to capture high-frequency features and improves overall stability. FEDONet achieves lower pointwise errors while eliminating the artificial spectral plateau that characterizes conventional fully connected trunks.
3.2 Burgers’ Equation
The Burgers’ equation is a well-known 1D viscous equation. It has been studied extensively for DeepONets in prior literature [53]. The equation is given by:
$\frac{\partial u}{\partial t}(x,t)+u\frac{\partial u}{\partial x}(x,t)=\nu\frac{\partial^{2}u}{\partial x^{2}}(x,t),\>\>\>\>\>\forall(x,t)\in[0,1]\times[0,1]$ (13)
where $x$ and $t$ denote the spatio-temporal coordinates and $\nu$ is the kinematic viscosity. We set $\nu=0.01$ along with initial and periodic boundary conditions as:
$\begin{split}\displaystyle u(x,0)=&\>s(x),\>\>\>\>\>\>\>\>\>\>\>\forall\>x\in[0,1]\\ u(0,t)=&\>u(1,t),\>\>\>\>\>\>\>\>\forall\>t\in[0,1]\\ \frac{\partial u}{\partial x}(0,t)=&\>\frac{\partial u}{\partial x}(1,t)\>\>\>\>\>\>\forall\>t\in[0,1]\end{split}$ (14)
where the initial condition $s(x)$ is sampled from a Gaussian Random Field, $s(x)\sim\mathcal{N}\left(0,25^{2}(-\Delta+5^{2}I)^{-4}\right)$, satisfying periodicity. The objective is to learn the nonlinear operator $\mathcal{G}$ mapping $s(x)$ to the full spatio-temporal solution $u(x,t)$ using both vanilla DeepONet and the proposed FEDONet.
Both models were evaluated on $250$ unseen test samples, yielding average relative $L^{2}$ errors of $6.43\%$ for the DeepONet and $4.86\%$ for the FEDONet. Figure 4 presents prediction for a test sample, showing the ground truth solution alongside predictions from both models. While DeepONet captures the coarse solution structure, it struggles with steep gradients and exhibits visible smoothing near shock regions. In contrast, FEDONet produces significantly sharper and more accurate reconstructions, faithfully capturing shock fronts and fine-scale dynamics.
Figure 4: Solving a parametric Burgers’ equation: Comparison of predicted spatio-temporal field for a random initial condition. Relative $L^{2}$ error: $4.09\%$ for DeepONet, $1.49\%$ for FEDONet.
The energy spectra of the predicted and ground truth fields, shown in Figure 5, further emphasize this advantage. FEDONet accurately captures the spectral decay across a wide range of wavenumbers, aligning with the physical expectations of dissipative transport phenomena. DeepONet, however, underestimates energy at higher modes, reflecting its inability to resolve fine features and spectral richness. This demonstrates that the Fourier embeddings effectively bridge the model’s spectral bias, enabling it to better represent high-frequency dynamics essential to Burgers-type flows. The enhanced performance of FEDONet in both physical and spectral domains highlights its potential as a more robust and generalizable operator learner for convection-dominated systems.
Figure 5: Energy spectrum and pointwise absolute error across space and time for a representative sample for Burgers’ Equation.
Figure 6 shows the convergence of training losses for different batch sizes. FEDONet consistently converges faster and to lower loss values across all training configurations, indicating improved optimization characteristics and stability during training.
Figure 6: Training loss convergence for the 1D Burgers’ Equation - Comparison of DeepONet and FEDONet trained with varying batch sizes.
3.3 Lorenz-63 System
The Lorenz-63 system is a widely studied chaotic dynamical system that serves as a benchmark for evaluating the long-term stability and predictive accuracy of surrogate models. It is governed by the following system of coupled nonlinear ordinary differential equations:
$\frac{dx}{dt}$ $=\sigma(y-x),$ (15)
$\frac{dy}{dt}$ $=x(\rho-z)-y,$
$\frac{dz}{dt}$ $=xy-\beta z,$
where the canonical parameters are set to $\sigma=10$, $\rho=28$, and $\beta=8/3$. Under these settings, the system exhibits chaotic dynamics on a strange attractor in three-dimensional phase space.
The objective is to approximate the solution operator $\mathcal{G}\colon x_{0}\mapsto\mathbf{x}(t)$, where $\mathbf{x}(t)=(x(t),y(t),z(t))$ denotes the full trajectory over the time interval $t\in[0,3]$, given a scalar initial condition $x_{0}\in[10,15]$, with $y_{0}=z_{0}=12$ held fixed. This configuration reduces the input to a one-dimensional function-to-function mapping, while preserving the rich, sensitive behavior of the output trajectory.
A dataset of 10,00010{,}000 trajectories is generated using the classic Runge-Kutta method (RK45), each evaluated at 1000 uniformly spaced time steps. The input to the DeepONet branch network is the scalar $x_{0}$, and the trunk network is queried at time coordinates $t\in[0,3]$. The goal is to accurately recover the full state trajectory, including lobe transitions and fine-scale oscillations characteristic of chaotic dynamics.
Figure 7: Reconstructed Lorenz-63 attractor for a randomnly sampled initial condition for trained DeepONet and FEDONet
Figure 7 shows the reconstructed attractor for a representative test sample. While DeepONet departs from the ground truth early in the rollout, FEDONet maintains geometric coherence and captures both wings of the attractor. This result demonstrates the stabilizing effect of spectral enrichment on trajectory learning.
Figure 8: Phase-plane projection of the xxzz states.
In Figure 8, we examine the xxzz projection. The DeepONet prediction deviates from the true phase-space trajectory and fails to capture the correct lobe transitions. In contrast, FEDONet successfully recovers the attractor’s phase geometry, indicating more robust learning of long-term dependencies.
Figure 9: Pointwise prediction error in $x(t)$, $y(t)$, and $z(t)$ for a representative sample.
Figure 9 presents the pointwise prediction errors over time. DeepONet accumulates noticeable errors, particularly in the $x(t)$ and $z(t)$ components, which undergo rapid transitions between lobes. These deviations grow sharply after $t>1.5$, illustrating compounding approximation errors. FEDONet exhibits substantially lower and smoother error profiles, with better robustness during nonlinear state transitions.
Figure 10: Cumulative relative error over time.
In Figure 10, we observe that the cumulative error of DeepONet grows exponentially, characteristic of chaotic divergence. In contrast, FEDONet maintains a flatter cumulative error curve, demonstrating enhanced resilience against long-term instability and better preservation of trajectory fidelity.
Figure 11: Spectral energy density of predicted $x(t)$ trajectories compared with the ground truth.
Finally, Figure 11 compares the spectral content of the predicted and reference trajectories. The vanilla DeepONet fails to capture high-frequency modes, resulting in a smoothed trajectory that diverges from the correct dynamics. FEDONet, equipped with randomized Fourier features, preserves the broadband energy spectrum more accurately - an essential factor in recovering the fine-scale structure and temporal richness of chaotic systems.
These results collectively demonstrate that FEDONet offers improved phase alignment, lower pointwise and cumulative error, and superior spectral accuracy in chaotic dynamical systems. The introduction of Fourier embeddings into the trunk network acts as a spectral preconditioner, enabling the model to overcome inductive biases and improve fidelity under sensitivity to initial conditions.
3.4 Eikonal Equation
We consider the two-dimensional Eikonal equation, which governs the propagation of wavefronts at unit speed and arises in a wide range of applications including geometric optics, level-set methods, and computational geometry. The equation is given by:
$\|\nabla s(\mathbf{x})\|_{2}$ $=1,$ (16)
$s(\mathbf{x})$ $=0,\quad\mathbf{x}\in\partial\Omega,$
where $\mathbf{x}=(x,y)\in\mathbb{R}^{2}$ denotes the spatial coordinates, $\Omega\subset\mathbb{R}^{2}$ is the computational domain, and $\partial\Omega$ is its boundary. The solution $s(\mathbf{x})$ represents the signed distance from a point $\mathbf{x}\in\Omega$ to the boundary $\partial\Omega$, defined as:
[…]
with the distance function $d$ defined by:
$d(\mathbf{x},\partial\Omega):=\inf_{\mathbf{y}\in\partial\Omega}\|\mathbf{x}-\mathbf{y}\|_{2}.$ (18)
In this study, we aim to learn the solution operator that maps a binary mask of a two-dimensional geometry to its corresponding signed distance field (SDF). As a testbed, we construct a dataset of NACA 4-digit airfoil shapes, widely used in aerodynamics, and compute their SDFs within a structured computational domain.
A NACA 4-digit airfoil is parameterized by its maximum camber $m$, the location of maximum camber $p$, and the maximum thickness $t$. The mean camber line is given by:
$y_{c}(x)=\begin{cases}\frac{m}{p^{2}}(2px-x^{2}),&0\leq x<p,\\ \frac{m}{(1-p)^{2}}((1-2p)+2px-x^{2}),&p\leq x\leq 1,\end{cases}$ (19)
and its derivative is:
$\frac{dy_{c}}{dx}=\begin{cases}\frac{2m}{p^{2}}(p-x),&0\leq x<p,\\ \frac{2m}{(1-p)^{2}}(p-x),&p\leq x\leq 1.\end{cases}$ (20)
The thickness distribution $y_{t}(x)$ is defined as:
$y_{t}(x)=5t\left(0.2969\sqrt{x}-0.1260x-0.3516x^{2}+0.2843x^{3}-0.1036x^{4}\right).$ (21)
The upper and lower surface coordinates of the airfoil are computed as:
$x_{u}=x-y_{t}\sin(\theta),\quad$ $y_{u}=y_{c}+y_{t}\cos(\theta),$ (22)
$x_{l}=x+y_{t}\sin(\theta),\quad$ $y_{l}=y_{c}-y_{t}\cos(\theta),$ (23)
where $\theta=\tan^{-1}\left(\frac{dy_{c}}{dx}\right)$ is the local inclination angle of the camber line.
Each airfoil is embedded in a $256\times 256$ pixel computational grid and scaled to fit within a central subdomain. A binary mask is created to define the geometry, and the signed distance function is computed using the Euclidean Distance Transform (EDT) as:
$s(x,y)=d_{\text{out}}(x,y)-d_{\text{in}}(x,y),$ (24)
where $d_{\text{out}}$ and $d_{\text{in}}$ represent the unsigned distances outside and inside the airfoil boundary, respectively. To ensure numerical stability, the resulting SDF is normalized to the range [1,1][-1,1]:
$s_{\text{norm}}(x,y)=\frac{s(x,y)}{\max|s(x,y)|}.$ (25)
A total of 1250 distinct airfoil shapes are synthesized by randomly sampling parameters $m\in[0.01,0.09]$, $p\in[0.1,0.7]$, and $t\in[0.1,0.4]$. Of these, 1000 samples are used for training and 250 for testing. This dataset captures a wide range of geometrical variations and curvature profiles, providing a rigorous benchmark for learning PDE solution operators from geometric input. On this benchmark, the baseline DeepONet attains an average relative $L^{2}$ error of 2.15%, while the proposed FEDONet achieves an improved 1.12% error, demonstrating substantially higher reconstruction accuracy.
Figure 12: Solving a parametric Eikonal equation (airfoils): Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry. Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
As shown in Figure 12, FEDONet achieves more accurate recovery of the level-set structure. The DeepONet underperforms in regions of high curvature, particularly near the airfoil’s leading and trailing edges, where geometric non-smoothness introduces steep spatial gradients. FEDONet’s use of Fourier embeddings in the trunk network enhances its ability to resolve these features by expanding the function space with oscillatory basis functions aligned with the data’s spectral demands.
These results confirm that FEDONet offers improved resolution of high-gradient and discontinuous regions in PDE fields, even in geometry-conditioned tasks such as the Eikonal equation. This highlights its potential as a general-purpose surrogate for PDEs defined on complex, spatially varying domains.
3.5 Allen–Cahn Equation
We further evaluate the performance of the proposed FEDONet in modeling the one-dimensional Allen–Cahn nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems. The governing partial differential equation is given by:
$\frac{\partial u}{\partial t}=\epsilon\frac{\partial^{2}u}{\partial x^{2}}-5u^{3}+5u,\quad x\in[-1,1],\quad t\in[0,1],$ (26)
where $u(x,t)$ is the phase field variable and $\epsilon=10^{-4}$ denotes the diffusion coefficient. Periodic boundary conditions are enforced on both $u$ and its spatial derivative $\partial u/\partial x$, consistent with the physical assumption of a closed domain.
To construct a challenging dataset with rich spatial features, we generate 10,000 solution trajectories using an explicit Euler solver. The spatial and temporal grids are discretized using $\Delta x=0.01$ and $\Delta t=0.005$, yielding solution arrays of dimension $200\times 200$. Initial conditions are defined as:
$s(x)=\sum_{k=1}^{3}\left[a_{k}x^{2k}\cos(k\pi x)+b_{k}x^{2k}\sin(k\pi x)\right],$ (27)
where $a_{k},b_{k}\sim\mathcal{U}(0,1)$ are independent samples from the uniform distribution. This construction introduces multiscale structure and localized steep gradients into the initial data, posing a nontrivial test for surrogate learning.
Each training sample comprises a pair $(s(x),u(x,t))$, where $s(x)\in\mathbb{R}^{200\times 1}$ denotes the initial profile, and $u(x,t)\in\mathbb{R}^{200\times 200}$ is the full spatiotemporal solution. The goal is to approximate the solution operator $\mathcal{G}\colon s(x)\mapsto u(x,t)$, comparing the performance of vanilla DeepONet and FEDONet.
Figure 13: Solving a parametric Allen Cahn equation: Comparison of predicted spatio-temporal field for a random initial condition. Relative $L^{2}$ error: $4.85\%$ for DeepONet, $2.27\%$ for FEDONet.
Both models were evaluated on $128$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $10.61\%$ for the DeepONet and $5.50\%$ for the FEDONet. Figure 13 shows qualitative predictions for a representative test case. FEDONet sharply resolves the interface zones and phase transitions, while DeepONet produces smoother approximations with notable blurring near discontinuities. The pointwise absolute error plot in Figure 14 emphasizes the localized improvements of FEDONet in capturing steep interfaces. DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
Figure 14: Energy spectrum and pointwise absolute error across space and time for a representative sample for Allen-Cahn Equation.
Figure 14 compares the spatial energy spectra of predicted and true solutions. FEDONet accurately tracks the ground truth spectrum across all wavenumbers, preserving the high-frequency tail associated with sharp phase boundaries. In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen-Cahn system.
3.6 Kuramoto–Sivashinsky Equation
We evaluate the performance of FEDONet on the one-dimensional Kuramoto-Sivashinsky (KS) equation, a canonical model that exhibits spatiotemporal chaos due to the combined effects of nonlinear advection, destabilizing diffusion, and stabilizing fourth-order hyperviscosity. The governing equation is given by:
$\begin{split}&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{4}u}{\partial x^{4}}=0,\quad x\in[0,L],\;t\in[0,T],\\ &u(x,0)=u_{0}(x),\quad x\in[0,L],\end{split}$ (28)
where $u(x,t)$ is the scalar field of interest, and $L=24$, $T=50$ denote the spatial and temporal domains respectively. The initial condition $u_{0}(x)$ is synthesized as a randomized Fourier series:
$u_{0}(x)=\sum_{n=1}^{4}C_{n}\sin\left(\frac{nx}{L}\right),\quad C_{n}\sim\mathcal{N}(0,1),$ (29)
which introduces low-frequency coherent structures that quickly evolve into chaotic waveforms under the KS dynamics.
We generate 10,000 such initial conditions and simulate the corresponding solution fields using a pseudo-spectral solver. The objective is to learn the operator $\mathcal{G}\colon u_{0}(x)\mapsto u(x,t)$ using both the standard DeepONet and the proposed FEDONet models. This problem serves as a stringent benchmark due to the sensitivity of chaotic systems to initial conditions and their broadband frequency content.
Figure 15: Solving a parametric KS equation: Median performing test-sample. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
Figure 15 shows the reconstructed spatiotemporal field for a median test case. The FEDONet model effectively captures the nonlinear evolution of the solution, preserving the wave envelope, phase dynamics, and intermittent modulation typical of KS attractors. In stark contrast, the vanilla DeepONet produces incoherent and smoothed reconstructions, with pronounced phase lag and energy collapse. This degradation is reflected in a high relative error of $89.63\%$ compared to just $16.36\%$ for FEDONet.
In Figure 16, we highlight the best-performing test case. FEDONet achieves a relative $L^{2}$ error of only $10.77\%$, accurately modeling both the fine-scale chaotic filaments and large-scale rolling structures. Notably, the trajectory remains consistent with the ground truth over long integration times, demonstrating that the spectral trunk can stabilize predictions and suppress drift. Such performance underscores the model’s ability to retain energy in high-frequency modes essential for modeling non-periodic fine-grained behavior that typifies chaotic fields.
Figure 16: Solving a parametric KS equation: Best performing test sample. FEDONet retains the amplitude and frequency content of the chaotic trajectory, yielding a low relative $L^{2}$ error of $10.77\%$.
Figure 17: Solving a parametric KS equation: Worst performing test sample. Even in failure cases, FEDONet remains qualitatively reasonable. Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
Figure 17 displays the worst-case test result. While the FEDONet prediction exhibits growing deviations from the reference field, it still retains discernible wave structures and phase coherence, with an error of $58.02\%$. In comparison, the vanilla DeepONet fails entirely, producing numerical artifacts, amplitude distortion, and full trajectory collapse. This robustness under extreme conditions emphasizes the value of frequency-enhanced representations in learning dynamical system operators.
Figure 18 plots the training loss curves for both models. FEDONet converges faster and exhibits more stable optimization behavior, attributable to the improved conditioning of the input space via randomized Fourier features. This aligns with observations from kernel learning and feature engineering literature, where frequency-based transformations are known to facilitate smoother descent landscapes and gradient propagation.
Figure 18: KS Training convergence comparison. FEDONet achieves faster and more stable optimization.
Figure 19: Distribution of relative $L^{2}$ errors across 512 unseen test samples for KS Equation. FEDONet maintains lower error in both mean and variance, demonstrating improved robustness and generalization in chaotic regimes.
Finally, Figure 19 summarizes performance statistics across 512 randomly selected test cases. FEDONet consistently achieves lower relative $L^{2}$ errors with reduced variance, indicating better generalization and resilience to initial condition perturbations.
In totality, the Kuramoto-Sivashinsky benchmark clearly illustrates the limitations of standard DeepONet trunks in handling chaotic, broadband dynamics. The introduction of Fourier embeddings into the trunk network provides an effective inductive bias that promotes spectral alignment, temporal coherence, and long-horizon accuracy. FEDONet emerges as a compelling model for operator learning in complex regimes, particularly those characterized by sensitive dependence on initial conditions and multiscale feature interactions.
4 Summary
In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations. By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training all without increasing runtime complexity. FEDONet consistently outperforms the vanilla DeepONet. Notably, it delivers dramatic reductions in relative $L^{2}$ error, most prominently in multiscale and high-frequency scenarios such as the Kuramoto-Sivashinsky equation (76% $\rightarrow$ 21%).
Table 1: Relative $L^{2}$ error (%) for each benchmark.
Dataset DeepONet FEDONet
2-D Poisson (elliptic) 1.41 1.09
1-D Burgers (conv./diff.) 6.43 4.86
Lorenz-63 (chaotic ODE) 2.03 0.46
Allen-Cahn (phase–field) 10.61 5.50
Kuramoto-Sivashinsky (spatio-chaos) 75.11 20.62
Eikonal (Hamilton–Jacobi) 2.15 1.12
Beyond pointwise accuracy, FEDONet preserves physical consistency across modalities. Energy spectrum analyses reveal that FEDONet better captures inertial and dissipation range behavior, avoiding the artificial spectral pile-up observed in DeepONets. In geometry-sensitive tasks like the Eikonal equation, it reconstructs sharp features and discontinuities with higher fidelity. Taken together, these results establish FEDONet as a robust, scalable, and architecture-agnostic enhancement for neural operator learning. The proposed spectral trunk serves as a drop-in replacement that aligns the inductive biases of DeepONets with the spectral demands of complex PDE systems, providing a promising path forward for efficient surrogate modeling in scientific machine learning.
5 Future Work
The Fourier-Embedded DeepONet (FEDONet) architecture presents a compelling foundation for enhancing operator learning, yet several promising extensions remain unexplored. A natural next step is to generalize beyond sinusoidal embeddings by investigating alternative basis functions in the trunk network. Wavelets, Chebyshev polynomials, or orthogonal polynomials tailored to specific PDE families could offer improved localization or boundary adaptivity, especially for systems exhibiting sharp interfaces or multi-scale dynamics. Additionally, hybridizing spectral embeddings with structured architectures such as graph-based trunks or encoder-decoder networks may further boost performance on irregular domains or hierarchical solution spaces.
Another avenue involves making the Fourier embedding parameters learnable. While the current design uses fixed random frequency matrices, enabling adaptive tuning of these frequencies could allow the model to better align with the intrinsic spectral content of the data. This modification could yield a more expressive and data-efficient hypothesis space while preserving the training stability benefits of approximate whitening. Finally, incorporating uncertainty quantification and physics-informed constraints into the FEDONet framework via Bayesian formulations or PINO style residual losses may further enhance robustness, interpretability, and generalization in scientific computing tasks involving sparse or noisy data.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This work was supported in part by the AFOSR Grant FA9550-24-1-0327.
Data availability
Data supporting the findings of this study are available from the corresponding author upon reasonable request.
Declaration of Generative AI use
During the preparation of this work, the authors used ChatGPT and Grammarly to improve readability and language. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Appendix A Whitening Effect of Fourier Feature Embeddings
Let $\phi(\zeta)=\sqrt{2}\,[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)]\in\mathbb{R}^{2M}$ be the Fourier feature embedding of an input $\zeta\in\mathbb{R}^{d}$, where each row of the matrix $Z\in\mathbb{R}^{M\times d}$ is sampled i.i.d. from a Gaussian distribution:
$Z_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (30)
Assume further that the input $\zeta$ is uniformly distributed over a compact domain, e.g., $\zeta\sim\mathcal{U}([0,1]^{d})$. Under these assumptions, we show that the random Fourier features $\phi(\zeta)$ are approximately whitened, i.e.,
$\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I_{2M}.$ (31)
To see this, consider a single frequency row vector $Z_{i}\in\mathbb{R}^{1\times d}$ drawn from $Z$. The corresponding $2$-dimensional block of the embedding is:
$\phi_{i}(\zeta)=\sqrt{2}\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)\\ \cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}$. (32)
We now compute the second-order statistics of this embedding block under the assumption that $Z_{i}\cdot\zeta$ is uniformly distributed over $[0,1]$ (a reasonable approximation due to the randomness in $Z_{i}$ and the uniformity of $\zeta$). Using standard trigonometric integrals, we obtain:
$\mathbb{E}_{\zeta}[\sin^{2}(2\pi Z_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (33)
$\mathbb{E}_{\zeta}[\cos^{2}(2\pi Z_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (34)
$\mathbb{E}_{\zeta}[\sin(2\pi Z_{i}\cdot\zeta)\cos(2\pi Z_{i}\cdot\zeta)]$ $\approx 0.$ (35)
Therefore, the unscaled covariance matrix for this block is approximately:
$\mathbb{E}_{\zeta}\left[\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)\\ \cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)&\cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}\right]\approx\frac{1}{2}I_{2}.$ (36)
Scaling by the $\sqrt{2}$ factor yields:
$\mathbb{E}_{\zeta}[\phi_{i}(\zeta)\phi_{i}(\zeta)^{\top}]\approx I_{2}.$ (37)
Since the rows of $Z$ are sampled independently and the embeddings for each $Z_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies:
$\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I_{2M}.$ (38)
Hence, the random Fourier features are approximately whitened. This whitening effect implies that the feature coordinates are uncorrelated and have unit variance. In practice, this leads to improved numerical conditioning and convergence in gradient-based optimization, particularly via the Neural Tangent Kernel (NTK), which benefits from isotropic input distributions. As such, Fourier embeddings not only enhance expressivity but also promote well-conditioned training dynamics.
Appendix B Theoretical Justification for Fourier Embedding Superset Property
In this appendix, we formally justify the claim stated in Section 2.1 that the hypothesis space induced by Fourier embeddings strictly contains the vanilla hypothesis space of coordinate-input networks. We present this as a lemma and provide constructive proof using classical function approximation theory.
Let $\mathcal{H}_{\text{vanilla}}$ be the hypothesis space of functions representable by a feedforward neural network with input $\zeta\in\mathbb{R}^{d}$, and let $\mathcal{H}_{\phi}$ be the hypothesis space of functions representable by a similar network where the input is first mapped via the Fourier feature embedding:
$\phi(\zeta)=\left[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)\right],\quad Z\in\mathbb{R}^{M\times d},\;Z_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (39)
Then, under mild regularity assumptions (e.g., compact input domain $K\subset\mathbb{R}^{d}$ and continuous activation functions), we have:
$\mathcal{H}_{\text{vanilla}}\subsetneq\mathcal{H}_{\phi}.$ (40)
We divide the proof into two parts.
Part I: Inclusion. Let $f\in\mathcal{H}_{\text{vanilla}}$ be a function representable by a neural network of the form:
$f(\zeta)=\sum_{j=1}^{N}a_{j}\,\sigma(w_{j}^{\top}\zeta+b_{j}).$ (41)
Since each $\zeta_{i}$ (coordinate function) can be approximated by a Fourier series on a compact domain $K$, and the Fourier embedding $\phi(\zeta)$ contains these sinusoidal components, it follows that a function of $\zeta$ can be emulated by a function of $\phi(\zeta)$. In particular, for every $f\in\mathcal{H}_{\text{vanilla}}$ and $\varepsilon>0$, there exists $\tilde{f}\in\mathcal{H}_{\phi}$ such that:
$\sup_{\zeta\in K}|f(\zeta)-\tilde{f}(\zeta)|<\varepsilon.$ (42)
Therefore, $\mathcal{H}_{\text{vanilla}}\subseteq\mathcal{H}_{\phi}$.
Part II: Strict Containment. Consider the function $f(\zeta)=\sin(100\pi\zeta)$ over $\zeta\in[0,1]$. Due to the spectral bias of neural networks [51], this high-frequency function is difficult to approximate using shallow networks on raw input. However, if the frequency matrix $Z$ used in the embedding $\phi(\zeta)$ includes rows close to $100$, then:
$\sin(2\pi Z_{j}\zeta)\approx\sin(100\pi\zeta),$
making $f$ easily representable in $\mathcal{H}_{\phi}$.
Since $f(\zeta)$ cannot be efficiently approximated in $\mathcal{H}_{\text{vanilla}}$ (requiring significantly more parameters or depth), but is trivially accessible in $\mathcal{H}_{\phi}$, we conclude that the inclusion is strict:
$\mathcal{H}_{\text{vanilla}}\subsetneq\mathcal{H}_{\phi}.$ (43)
This result formalizes the intuition that Fourier embeddings expand the expressivity of the network and can be interpreted as a form of spectral preconditioning for operator learning models such as DeepONets.
Appendix C Spectral Interpretation of Fourier-Embedded DeepONets
The use of Fourier embeddings in the trunk network induces a natural spectral decomposition in the DeepONet output. Specifically, the composition of the trunk and branch networks can be interpreted as a data-driven spectral projection:
$\mathcal{G}(u)(\zeta)\approx\sum_{k=1}^{p}\langle u,\varphi_{k}\rangle\,\psi_{k}(\zeta),$ (44)
where $\{\varphi_{k}\}_{k=1}^{p}$ are data-dependent basis functions learned from the branch network, and $\{\psi_{k}(\zeta)\}_{k=1}^{p}$ are basis functions defined by the trunk network, shaped by the sinusoidal feature embedding $\phi(\zeta)$.
When the trunk network is fed sinusoidal embeddings of the form
$\phi(\zeta)=[\cos(2\pi\omega_{1}^{\top}\zeta),\dots,\cos(2\pi\omega_{M}^{\top}\zeta),\sin(2\pi\omega_{1}^{\top}\zeta),\dots,\sin(2\pi\omega_{M}^{\top}\zeta)],$
the basis functions $\psi_{k}(\zeta)$ are learned combinations of trigonometric modes. This allows the model to represent oscillatory and localized patterns that are characteristic of multiscale physical systems.
The representation in Equation (44) mirrors classical Galerkin methods, where the solution is projected onto a known orthonormal basis:
$u(x,t)\approx\sum_{n=1}^{N}a_{n}(t)\,\phi_{n}(x),$ (45)
with the coefficients $a_{n}(t)$ governed by PDE dynamics. In DeepONet, both the basis functions $\psi_{k}$ and the projection coefficients $\langle u,\varphi_{k}\rangle$ are learned directly from data, enabling generalization beyond handcrafted bases.
Unlike classical global bases such as Fourier or Chebyshev functions, the neural trunk network can generate localized, adaptive basis functions. This enhances the model’s ability to approximate functions with sharp transitions, discontinuities, or compactly supported features, which are common in real-world PDE systems. We may formalize the DeepONet architecture as a composition of a projection and synthesis operator:
$\Pi_{\varphi}(u):=[\langle u,\varphi_{1}\rangle,\dots,\langle u,\varphi_{p}\rangle]^{\top},\quad\mathcal{S}_{\psi}(\zeta):=[\psi_{1}(\zeta),\dots,\psi_{p}(\zeta)]^{\top},$
yielding the output:
$\mathcal{G}(u)(\zeta)\approx\mathcal{S}_{\psi}(\zeta)^{\top}\Pi_{\varphi}(u),$
which resembles a spectral synthesis operation of the form:
$f(x)\approx\sum_{k}\hat{f}_{k}\phi_{k}(x).$
This interpretation leads to several insights:
  1. 1.
    Fourier-Embedded DeepONets naturally align with the dominant spectral modes of the target operator, facilitating more efficient approximation of high-frequency content.
  2. 2.
    The learned basis can adapt to complex patterns in the data, unlike fixed basis expansions.
  3. 3.
    The operator learning framework parallels well-established numerical discretization techniques, such as Petrov-Galerkin and spectral methods, but operates in a fully data-driven regime.
In summary, Fourier Embeddings serve as a principled mechanism for spectral lifting, enabling DeepONet to perform a learned Galerkin-style decomposition of operators. This perspective bridges classical PDE approximation theory with modern deep learning, offering both interpretability and flexibility in learning complex function-to-function mappings.